# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v4

@@ Acknowledgments
We consider remote sensing data  (?, ?) with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
$R_{t}\in\mathbb{R}^{W\times H\times C}$ (S1)
@@
$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (S3)
$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (S4)
@@
We represent each 1- m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig. 1 a). In this array, each row corresponds to a chronologically ordered observation date, and each column is a specific spectral band or polarization. This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities. Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored in the temporal sampling process.
@@
Each encoder begins by linearly embedding the input features (spectral bands or polarizations) for each time step. To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings. The core of each encoder consists of a stack of 4 standard Transformer blocks  (?), featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
@@
The fused embedding from the dual-encoder stage is subsequently fed into a large projector network. This projector is a six-layer MLP. Its architecture comprises an input layer mapping the fused embeddings to 16,384 dimensions, followed by four hidden layers of 16,384 dimensions each, and a final linear output layer. Each of the first five layers is a fully-connected linear layer followed by Batch Normalization and a ReLU non-linear activation function, making the network deep and highly non-linear. This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work  (?), to enable effective redundancy reduction during the self-supervised loss computation. The final output of the projector is then used in the loss calculation. For downstream tasks, we used the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding. The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
@@
The TESSERA model is trained using a modified Barlow Twins objective function  (?). For this objective, two distorted views, denoted as $Y_{A}$ and $Y_{B}$, are generated for each input d-pixel. In TESSERA, these views are created by independently running the temporal sampling and preprocessing pipeline twice for the Sentinel-1 and Sentinel-2 data associated with a given d-pixel. This process involves:
@@
The network processes these two views ($Y_{A},Y_{B}$) through the dual-encoder and the projector to produce batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as  (?):
$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (S7)
Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The indices $i$ and $j$ iterate over the dimensions of the embedding vectors. The first term (invariance term) encourages similar embeddings for different views of the same input ($C_{ii}\to 1$). The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$. We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference  (?).
To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al.  (?). For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$. The mixing coefficient $\alpha_{mix}$ is sampled from a uniform distribution on the unit interval, $\alpha_{mix}\sim U(0,1)$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained from their respective views. The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (S8)
$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (S9)
$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (S10)
@@
$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (S11)
where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by  (?)) to be effective based on our experiments.
The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles. d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400. This was achieved by systematically taking every 20th pixel along the horizontal and vertical axes. Each d-pixel contained the annual time series for 10 spectral bands of Sentinel-2 data and the annual time series of VV and VH polarizations for Sentinel-1 data.
@@
As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training. The global batch size was 32768. Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary fig. S2.
A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset. Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch. To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2 TB of initial d-pixel data. This pipeline is conceptually illustrated in Fig. S3a.
The impact of this approach on training stability is empirically demonstrated in Fig. S3b. Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
@@
A primary output of the TESSERA project is the generation of annual global embedding maps with 1- meter resolution for the years 2017-2024. To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used. For each 1- meter pixel on the globe and for each year:
@@
    The full Sentinel-1 and Sentinel-2 time series data at 1- meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
@@
This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 1- meter grid.
Scaling with data and network size
To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task. Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig. S4).
@@
In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning. As detailed in Section Projector Network, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers. Our parameter sweep focused on this computationally intensive hidden block. We varied the number of these hidden layers (referred to as ’depth’ in Fig. S4a) and their width (number of neurons per layer), with results shown in Fig. S4a. As discussed by Zbontar et al.  (?), deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task. The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide). However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure. For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig. S4b. The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget. This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks. The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
@@
Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score  (?). The RankMe score quantifies the richness of the embeddings by measuring how uniformly the information is distributed across the feature dimensions. Given a batch of embeddings represented by the matrix $Z\in\mathbb{R}^{N\times D}$, where $N$ is the batch size and $D$ is the embedding dimension, we first compute its singular values $s_{1},s_{2},\ldots,s_{D}$. These are normalized to form a probability distribution $p_{i}=s_{i}/\sum_{j}s_{j}$. The RankMe score is the Shannon entropy of this distribution, normalized by the maximum possible entropy:
$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (S12)
@@
The results of the ablation study are summarized in Table S4. Our analysis reveals several key insights into the model’s architecture and training strategy. The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches. Removing mixup regularization also severely degrades results. When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
@@
Downstream Task Application Methodology

### v5

@@ A.1 Model details
We consider remote sensing data [100, 99] with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
$R_{t}\in\mathbb{R}^{W\times H\times C}$ (1)
@@
$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (3)
$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (4)
@@
We represent each 10 m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig. 1 a). In this array, each row corresponds to a chronologically ordered observation date, and each column is a specific spectral band or polarization. This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities. Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored in the temporal sampling process.
@@
Each encoder begins by linearly embedding the input features (spectral bands or polarizations) for each time step. To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings. The core of each encoder consists of a stack of 4 standard Transformer blocks [107], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
@@
The fused embedding from the dual-encoder stage is subsequently fed into a large projector network. This projector is a six-layer MLP. Its architecture comprises an input layer mapping the fused embeddings to 16,384 dimensions, followed by four hidden layers of 16,384 dimensions each, and a final linear output layer. Each of the first five layers is a fully-connected linear layer followed by Batch Normalization and a ReLU non-linear activation function, making the network deep and highly non-linear. This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [34], to enable effective redundancy reduction during the self-supervised loss computation. The final output of the projector is then used in the loss calculation. For downstream tasks, we used the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding. The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
Table 2: Architectural parameters of the TESSERA v1.0 model.
Component Parameters Percentage of Total
Encoder (Feature Extractor) 45,697,026 3.29%
   Sentinel-1 Encoder 22,846,465 1.64%
   Sentinel-2 Encoder 22,850,561 1.64%
Fusion MLP 131,200 <<0.01%
Projector (for Training only) 1,344,700,421 96.71%
Total Parameters 1,390,528,647 100.00%
@@
The TESSERA model is trained using a modified Barlow Twins objective function [34]. For this objective, two distorted views, denoted as $Y_{A}$ and $Y_{B}$, are generated for each input d-pixel. In TESSERA, these views are created by independently running the temporal sampling and preprocessing pipeline twice for the Sentinel-1 and Sentinel-2 data associated with a given d-pixel. This process involves:
@@
The network processes these two views ($Y_{A},Y_{B}$) through the dual-encoder and the projector to produce batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as [34]:
$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (7)
Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The indices $i$ and $j$ iterate over the dimensions of the embedding vectors. The first term (invariance term) encourages similar embeddings for different views of the same input ($C_{ii}\to 1$). The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$. We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference [34].
To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al. [108]. For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$. The mixing coefficient $\alpha_{mix}$ is sampled from a uniform distribution on the unit interval, $\alpha_{mix}\sim U(0,1)$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained from their respective views. The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (8)
$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (9)
$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (10)
@@
$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (11)
where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by [108]) to be effective based on our experiments.
The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles chosen from the years 2017-2024. d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400. This was achieved by systematically taking every 20th pixel along the horizontal and vertical axes. Each d-pixel contained the annual time series for 10 spectral bands of Sentinel-2 data and the annual time series of VV and VH polarizations for Sentinel-1 data.
@@
Figure 8: TESSERA is trained on globally-distributed training data. TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024. The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend. This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training. The global batch size was 32,768. Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary fig. 9.
A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset. Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch. To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2 TB of initial d-pixel data. This pipeline is conceptually illustrated in Fig. 10a.
The impact of this approach on training stability is empirically demonstrated in Fig. 10b. Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
Figure 9: Training dynamics curves show that TESSERA rapidly reaches its asymptotes. a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss. b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay. c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction. d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant. e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
@@
Figure 10: Data shuffling greatly improves training stability. a, Schematic of the data shuffling and loading process. D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary [37] performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
@@
A primary output of the TESSERA project is the generation of annual global embedding maps with 10 meter resolution for the years 2017-2024. To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used. For each 10 meter pixel on the globe and for each year:
@@
    The full Sentinel-1 and Sentinel-2 time series data at 10 meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
@@
This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10 meter grid.
A.2 Scaling with data and network size
To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task. Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig. 11).
@@ A.2 Scaling with data and network size
In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning. As detailed in Section A.1, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers. Our parameter sweep focused on this computationally intensive hidden block. We varied the number of these hidden layers (referred to as ’depth’ in Fig. 11a) and their width (number of neurons per layer), with results shown in Fig. 11a. As discussed by Zbontar et al. [34], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task. The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide). However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 16,384 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure. For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
Figure 11: Model performance improves with increasing projector size and additional computation. a, Hyperparameter sweep for the projector network’s architecture. Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis). The colour and size of the point indicate the validation F1 score and the number of parameters, respectively. Performance generally increases with both width and depth. The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints. The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency. b, Scaling laws for TESSERA. The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters). A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used. The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig. 11b. The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget. This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks. The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
Figure 12: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications. Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks. Users can then fine-tune the entire model or only the head. Another approach (b) is where both the encoder and the projector (or decoder) are typically updated. In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input. This results in lightweight downstream models that consist only of a task-specific head.
@@
Table 3: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance. We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset. The baseline represents the full model. Each subsequent row shows the result of removing the specified component. Performance is measured by validation F1 score and the RankMe score. Arrows ($\uparrow$) indicate that higher values are better. The results demonstrate that all components contribute positively to the model’s overall performance.
Baseline (Full Model) 0.773 0.963
w/o Global Data Shuffling 0.681 0.847
w/o Mixup Regularization 0.662 0.857
w/o Sentinel-1 Data 0.742 0.931
w/o Shuffling & Mixup 0.626 0.867
w/o Embedding Quantization 0.779 0.972
@@
Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score [109]. The RankMe score quantifies the richness of the embeddings by measuring how uniformly the information is distributed across the feature dimensions. Given a batch of embeddings represented by the matrix $Z\in\mathbb{R}^{N\times D}$, where $N$ is the batch size and $D$ is the embedding dimension, we first compute its singular values $s_{1},s_{2},\ldots,s_{D}$. These are normalized to form a probability distribution $p_{i}=s_{i}/\sum_{j}s_{j}$. The RankMe score is the Shannon entropy of this distribution, normalized by the maximum possible entropy:
$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (12)
@@
The results of the ablation study are summarized in Table 3. Our analysis reveals several key insights into the model’s architecture and training strategy. The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches. Removing mixup regularization also severely degrades results. When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
@@
A.3 Downstream Task Application Methodology